### v2

# Holistic Evaluation for Interleaved Text-and-Image Generation

Abstract
Interleaved text-and-image generation has been an intriguing research direction, where the models are required to generate both images and text pieces in an arbitrary order. Despite the emerging advancements in interleaved generation, the progress in its evaluation still significantly lags behind. Existing evaluation benchmarks do not support arbitrarily interleaved images and text for both inputs and outputs, and they only cover a limited number of domains and use cases. Also, current works predominantly use similarity-based metrics which fall short in assessing the quality in open-ended scenarios. To this end, we introduce InterleavedBench, the first benchmark carefully curated for the evaluation of interleaved text-and-image generation. InterleavedBench features a rich array of tasks to cover diverse real-world use cases. In addition, we present InterleavedEval, a strong reference-free metric powered by GPT-4o to deliver accurate and explainable evaluation. We carefully define five essential evaluation aspects for InterleavedEval, including text quality, perceptual quality, image coherence, text-image coherence, and helpfulness, to ensure a comprehensive and fine-grained assessment. Through extensive experiments and rigorous human evaluation, we show that our benchmark and metric can effectively evaluate the existing models with a strong correlation with human judgments surpassing previous reference-based metrics. We also provide substantial findings and insights to foster future research in interleaved generation and its evaluation. [^1] The source code and datasets are publicly available at https://huggingface.co/mqliu/InterleavedBench for research purposes.
Holistic Evaluation for Interleaved Text-and-Image Generation
Minqian Liu  Zhiyang Xu  Zihao Lin  Trevor Ashby  Joy Rimchala Jiaxin ZhangLifu Huang Virginia Tech  Intuit AI Research {minqianliu, zhiyangx, zihaol, trevorashby, lifuh}@vt.edu {joy_rimchala, jiaxin_zhang}@intuit.com
https://vt-nlp.github.io/InterleavedEval/
Figure 1: Overview of our InterleavedBench, a comprehensive benchmark that covers 10 diverse use cases for interleaved text-and-image generation, and the evaluation results of InterleavedEval based on GPT-4o.
1 Introduction
Figure 2: Comparison between the existing benchmark (multi-concept image composition Kumari et al. (2023a)) and our InterleavedBench. Compared with the existing benchmark, InterleavedBench has the following features: (1) both input and output can have arbitrarily interleaved text and images, and (2) each instance has a detailed instruction to benchmark models’ instruction-following capability.
Multimodal learning has been a rapidly developing research field given the recent advancements in Large Multimodal Models (LMMs) Xu et al. (2023); Dai et al. (2023); Liu et al. (2023a). While these models can perform diverse tasks such as detailed image description and visual question answering, the outputs are limited to the text-only format, which hinders their broader applications. More recently, there has been a growing focus on enhancing LMMs with the capability of interleaved generation, i.e., generating multimodal content that seamlessly integrates both text and one or multiple images Koh et al. (2023); Dong et al. (2024); Sun et al. (2023b, a). This opens new avenues for applications in diverse challenging scenarios, such as creative content generation Anantrasirichai and Bull (2022), visual storytelling Huang et al. (2016); Lukin et al. (2018), and multimodal script generation Yang et al. (2021); Qi et al. (2024).
While the LMMs for interleaved generation are continuously gaining stronger capabilities, progress in the evaluation of interleaved generation significantly lags behind with several critical challenges remaining. First, most existing works for interleaved generation quantitatively benchmark the models on text-to-image tasks where the output is usually one single image Koh et al. (2023); Dong et al. (2024). However, such evaluation methods would fail to assess model performance in the real-world scenarios of interleaved generation, where the output usually consists of interleaved text and images. Second, apart from human evaluation which is costly and time-consuming, existing works still heavily rely on reference-based metrics such as BLEU Papineni et al. (2002) FID Heusel et al. (2017) that measure the similarity between generated samples and gold references. Such similarity-based metrics often fail to accurately capture outputs’ quality, especially in open-ended tasks such as creative generation and visual storytelling. Third, the evaluation of interleaved generation is complex and involves many different aspects, such as perceptual quality, coherence between text and images, and helpfulness of the overall content. One single aspect is usually insufficient to reflect the overall quality. For example, despite the images in one output having good perceptual quality, the output can still be not helpful to users if the generated content is not coherent with the context, e.g., the request from users.
To address these critical limitations, we introduce InterleavedBench, the first benchmark for holistic evaluation of interleaved text-and-image generation. We construct InterleavedBench with a high-quality and diverse collection of interleaved generation scenarios that encompass a wide range of real-world use cases, including creative generation, multimodal script generation, visual storytelling, and many others. We compare our InterleavedBench and one existing benchmark Kumari et al. (2023b) closest to our dataset in Figure 2. To support the evaluation, we also introduce InterleavedEval, a strong reference-free evaluation metric based on GPT-4o OpenAI (2024), the current state-of-the-art LMM. InterleavedEval can take in any evaluation instructions and provide a fine-grained evaluation along with detailed explanations. We carefully curate a multi-aspect evaluation criterion to ensure a holistic evaluation for InterleavedEval. Specifically, we define five essential aspects for interleaved evaluation, including text quality, perceptual quality, image coherence, text-image coherence, and helpfulness, following the principles that (1) these aspects are generally applicable in different scenarios, (2) these aspects are atomic and orthogonal to each other, and (3) the combination of these aspects can comprehensively cover the critical dimensions in interleaved generation.
Extensive experiments and rigorous human evaluation demonstrate that (1) Our curated InterleavedBench posts unique and significant challenges to the existing integrated LMMs (e.g., GILL Koh et al. (2023) and EMU-2 Sun et al. (2023a)) for interleaved generation, where the quality of their outputs are far from satisfying. The pipeline systems combined with a strong LMM (e.g., GPT-4o) and a separate image generation model (e.g., DALLE3 Betker et al. ) generally achieve better results but still struggle on certain tasks; (2) InterleavedEval can achieve a good correlation with human judgments with significant improvement over previous automatic evaluation metrics; (3) The evaluation of interleaved generation remains a very challenging direction due to its complexity and the limitation of the existing LMM-based evaluator. We believe that our work can provide useful resources and insights for interleaved generation and its evaluation.
2 Related Work
Figure 3: Illustration of examples in our InterleavedBench from six representative use cases.
Large Multimodal Models for Interleaved Generation
The advent of large multimodal models (LMMs) Koh et al. (2023); Sun et al. (2023a) has significantly advanced the field of interleaved text-and-image generation. Previous models such as DALL-E Ramesh et al. (2021) and Stable Diffusion Podell et al. (2023) have demonstrated impressive capabilities in generating high-quality images conditioned on textual descriptions. However, previous focus has predominantly been on unidirectional generation tasks, either from text to image or image to text, without considering the interleaved generation scenarios where text and images are seamlessly integrated within the same output. Recent works have begun to address this gap, with the LMMs extended with diffusion models, exploring the generation of mixed text and image outputs Koh et al. (2023); Sun et al. (2023b); Dong et al. (2024); Tian et al. (2024); Zhan et al. (2024); Chen et al. (2023). These models leverage advanced architectures and training techniques to enhance their ability to produce coherent and contextually relevant interleaved content. Despite these advancements, the evaluation of such models remains an underexplored area, with most evaluations still relying on separate assessments of text and image quality or simplistic reference-based metrics. Our proposed InterleavedBench benchmark aims to bridge this gap by providing a holistic evaluation framework tailored specifically for interleaved text-and-image generation.
Evaluation of Multimodal Generation Tasks
Evaluating multimodal generation tasks presents unique challenges due to the inherent complexity of assessing both textual and visual components simultaneously. Existing metrics for text generation, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and LLM-based evaluators Zhong et al. (2022); Liu et al. (2023b, 2024), fall short when applied to multimodal outputs as they fail to capture the visual quality and coherence with textual content. Similarly, visual generation metrics like FID Heusel et al. (2017) and IS Salimans et al. (2016) are inadequate for evaluating the textual elements accompanying the images. To address this, recent studies have employed multimodal metrics Zhang et al. (2023b); Ku et al. (2023), such as CLIPScore Hessel et al. (2021), which leverages the alignment capabilities of the CLIP model to measure the similarity between generated images and their corresponding textual descriptions. However, CLIPScore can only measure the alignment between text and images, which is not sufficient to evaluate the quality of generated output comprehensively. Moreover, human evaluations, although more reliable, are resource-intensive and cannot be scalable. In terms of evaluation benchmarks in multimodal learning, existing works mostly focus on evaluating the tasks with single-modality output Fu et al. (2024); Li et al. (2024); Lu et al. (2024); Wang et al. (2024), such as conditional text-to-image generation Chen et al. (2024); Ku et al. (2024), where the primary focus is solely the quality of generated images. Our InterleavedBench benchmark introduces a novel approach to evaluate interleaved text-and-image generation by incorporating multiple aspects of quality assessment, thus providing a more nuanced and holistic evaluation framework.
3 InterleavedBench
We introduce InterleavedBench, the first comprehensive benchmark meticulously constructed to evaluate text-and-image interleaved generation. Figure 3 shows some examples from InterleavedBench.
3.1 Dataset Curation Process
Our dataset includes two subsets: a context-based subset where the instances contain a multimodal context of interleaved text and images in the input (first row in Figure 3), and a context-free subset with text-only inputs (second row in Figure 3). The context-free subset can assess whether the model can creatively generate interleaved content based on the text-only instruction, while the context-based subset can better benchmark the coherence and consistency of generated outputs.
Collection of Context-based Subset
Firstly, we collect the source data of the context-based subset from existing academic datasets or web resources. Specifically, we collect the data of multimodal script generation from WikiHow Yang et al. (2021), visual story completion from VIST Huang et al. (2016), activity generation from the dense captions and the extracted video frames in ActivityNet Captions Krishna et al. (2017), sequential image editing from MagicBrush Zhang et al. (2023a), and multi-concept image composition from CustomDiffusion Kumari et al. (2023a). For web resources, we apply an automatic data filtering pipeline to discard the samples with poor quality to obtain a small set of source data. We detail our data filtering pipeline in Appendix A. Secondly, after collecting the source data (either from academic benchmarks or web resources), we then apply a human selection process to manually select the samples based on data quality and diversity (i.e., avoiding selecting similar samples). Finally, we ask human experts to annotate an instruction $I$ for each sample based on the collected content. We include the details of the data selection and instruction annotation process in Appendix A. For the samples that are originally interleaved articles, we pick the first $k$ images and their associated text as the context $\mathcal{C}$ for the input. $k$ is randomly sampled for each example and ranges from 1 to the maximum number of images minus 1 since we need to ensure the output contains at least one image. The rest of the images and text are used as the gold reference.
Dataset Name Detailed Instruction Image Input Text Output Image Output
MagicBrush Zhang et al. (2023a) No Single No Single
DreamBench Chen et al. (2024) No Multiple No Single
CustomDiffusion Kumari et al. (2023a) No Multiple No Single
DreamEditBench Li et al. (2023) No Multiple No Single
Mantis-Eval Jiang et al. (2024) Yes Multiple Yes No
BLINK Fu et al. (2024) Yes Multiple Yes No
MuriBench Wang et al. (2024) Yes Multiple Yes No
InterleavedBench (Ours) Yes Multiple Yes Multiple
Table 1: Comparisons between InterleavedBench and existing open-sourced multimodal evaluation benchmarks. The highlighted features of our benchmark include detailed instructions and multiple images in input and/or output that are arbitrarily interleaved with text.
Figure 4: The distribution of the use cases in InterleavedBench.
Collection of Context-free Subset
The context-free subset consists of the use cases of marketing material generation, report generation, education content generation, and fairytale generation as they are common and practical scenarios for interleaved generation. We first leverage GPT-4o to generate a set of instances for each use case. For example, in marketing material generation, one instance is “creating marketing campaigns around holidays to boost sales”. Then, we use GPT-4o to extend each instance into a more detailed instruction, e.g., “Create an interleaved content that combines engaging text and eye-catching images for marketing campaigns around holidays to boost sales. Begin by researching holiday themes relevant to your products…”. Finally, we ask human annotators to verify whether the instructions are reasonable and of good quality. Note that we do not have gold references in this subset.
Dataset Statistics
In total, we finally collect 815 instances across 10 use cases, including multimodal script generation, document completion, visual story completion, marketing material generation, report generation, education content generation, activity generation, sequential image editing, and multi-concept image composition. The detailed distribution of the use cases is shown in Figure 4.
3.2 Comparison with Existing Benchmark
We highlight the following key differences and unique challenges introduced by our InterleavedBench compared with the existing benchmark. (1): Output modality: our benchmark requires the models to generate interleaved text and multiple images that could present in an arbitrary order, whereas exiting benchmarks Kumari et al. (2023b) only cover the output with single modality or a single image (as shown in Figure 2); (2) Requirement on coherence: given that both inputs and outputs in our benchmark can contain multiple pieces of text and images, our dataset can assess whether the outputs are coherent and consistent with input instruction and context, and within the outputs themselves; (3) Instruction following: Most existing conditional image generation datasets only contain simple instructions such as “add a cat next to the person”. On the contrary, each instance in our benchmark contains a detailed human-annotated instruction to describe the task. Thus, our dataset can evaluate models’ instruction-following and generalization capabilities. We show the difference between our benchmark and existing datasets in Table 1.
4 InterleavedEval
In many use cases of interleaved generation, such as “generate a story about Snow White using both text and images”, comparing the output against a gold reference is unrealistic since the generation can be fairly open-ended. However, existing approaches predominantly use reference-based metrics, e.g., BLEU Papineni et al. (2002) and FID Heusel et al. (2017), to measure the quality of text and image, respectively. They usually fail to assess the quality accurately.
To bridge the gap between existing metrics and the demand in more diverse and realistic scenarios, we present InterleavedEval, a strong reference-free metric based on GPT-4o, the current state-of-the-art LMM that supports arbitrarily interleaved inputs. To obtain a holistic and comprehensive evaluation of interleaved generation, we define five fine-grained evaluation aspects, including text quality, perceptual quality, image coherence, text-image coherence and helpfulness, and evaluate the output of each aspect separately. We show the detailed definition for each evaluation aspect in Table 5 in Appendix B. For each instance to be evaluated, the input of the evaluator consists of an instruction $I$ that indicates what should be accomplished, system output $X=(T_{O},\mathcal{P}_{O})$, where $T_{O}$ is the output text and $\mathcal{P}_{O}$ is the set of output images, the evaluation aspect $a$, and optionally, the context $\mathcal{C}$ of the task (e.g., the given text and images in models’ inputs).
We formulate the evaluation metric InterleavedEval as follows: We instruct the GPT-4o evaluator to output discrete scores from {0, 1, 2, 3, 4, 5} based on the detailed criteria shown in Table 5, where 1 indicates the worst quality, 5 indicates the best quality, and 0 indicates output text and/or images are empty. We also instruct GPT-4o to provide a detailed explanation to improve the interpretability. Note that when the output text is empty, the scores on text-related aspects (text quality and text-image quality) are 0. Similarly, when the output image is empty, the scores on image-related aspects (perceptual quality, image coherence, and text-image quality) are 0. Moreover, we do not apply the text-related aspects in sequential editing and subject-driven generation since the primary focus of these tasks is whether the image is generated correctly according to the instructions.
5 Experiments
Model Text Quality Perceptual Quality Image Coherence TIC Helpfulness AVG
MiniGPT-5 1.22 2.45 1.62 2.03 1.77 1.82
GILL 0.75 3.21 2.25 1.53 1.48 1.84
EMU-2 1.26 2.28 1.89 1.34 1.64 1.68
EMU-2 (Gold Text) 1.56 3.35 2.89 1.43 2.10 2.27
Gemini1.5 + SDXL 4.40 3.99 3.64 4.13 3.62 3.96
GPT-4o + DALL·E 3 4.37 4.36 3.51 4.55 3.88 4.13
Table 2: Automatic evaluation results of existing interleaved generation models on InterleavedBench using InterleavedEval. TIC is the abbreviation for ’Text-Image Coherence’. The best results are highlighted in bold.
Model Text Quality Perceptual Quality Image Coherence TIC Helpfulness AVG
GILL 1.35 1.89 1.72 1.43 1.19 1.52
EMU-2 1.23 1.74 1.87 1.24 1.2 1.46
Gemini1.5 + SDXL 2.59 2.36 2.13 2.27 2.08 2.28
GPT-4o + DALL·E 3 2.49 2.51 2.02 2.31 2.13 2.29
Table 3: Human evaluation results of existing interleaved generation models on InterleavedBench. TIC is the abbreviation for ’Text-Image Coherence’. The best results are highlighted in bold. Note that we use a scale of 0 to 3 for this evaluation, which is different from the scale used in Table 2.
Metric Ref-free? Text Quality Perceptual Quality Image Coherence TIC Helpfulness
BERTScore 0.21 - - - 0.37
DreamSim - 0.02 0.1 - 0.06
Image-Image CLIPScore - 0.08 0.2 - -0.01
Text-Image CLIPScore - - - 0.2 0.09
InterleavedEval 0.72 0.30 0.43 0.4 0.57
Table 4: Mete-evaluation on evaluation metrics in terms of Spearman correlation between automatic evaluation results with human judgments. For baseline metrics, we only report the correlation on the corresponding aspects (e.g., BERTScore can correspond to text quality) as well as helpfulness.
5.1 Experiment Setup
Baseline Models
We benchmark the following baseline models which can be categorized into two types: integrated models where the LMM and image generation model are connected via neural modules, and pipeline models where the LMM and image generation model are connected via prompts in natural language. The integrated models include: (1) MiniGPT-5 Zheng et al. (2023a) which connects a large language model with a stable diffusion model via generative vokens, enabling description-free multimodal generation; (2) GILL Koh et al. (2023) which allows a pretrained large language model to generate multimodal responses by mapping the hidden states of text into the embedding space of an image generation model; (3) EMU-2 Sun et al. (2023a) which induces in-context learning capabilities of LLMs by scaling up the model size and the size of the pretraining dataset; (4) EMU-2 Gen + Gold Text where EMU-2 Gen is a pretrained EMU-2 model instruction-tuned on various controllable image generation tasks. However, EMU-2 Gen cannot generate text so we combine it with ground-truth textual responses to come up with a complete text-and-image interleaved content for evaluation. The pipeline models include: (5) GPT-4o OpenAI (2024) + DALL·E 3 Betker et al. where GPT-4o is the state-of-the-art proprietary LMM that can comprehend interleaved text-and-image inputs and generate text-only responses. We leverage GPT-4o to generate text responses as well as captions for image responses in the desired positions. Then the captions are fed into DALL·E 3 to generate images. Finally, we combine the text responses with generated images in their original orders; (6) Gemini-1.5 Anil et al. (2023) + SDXL Podell et al. (2023): we build this baseline in a similar way as GPT-4o + DALL·E 3 but use Gemini-1.5 Pro as the LMM and Stable Diffusion XL Turbo as the image generation model.
Baseline Metrics
We adopt the following metrics as baselines to validate the effectiveness of our InterleavedEval. (1) BERTScore is a reference-based metric for text evaluation. We apply BERTScore to compute the similarity between the text output and the reference in our dataset. We set the BERTScore to 0 if the text output is empty. (2) CLIPScore is originally a reference-free evaluation metric for image captioning, which computes the cosine similarity between the CLIP embeddings of a predicted caption and that of the input image. We adopt CLIPScore as two baselines: a reference-based metric to compute image-image similarity between predicted images and ground truth images in a pair-wise manner, and a reference-free metric to compute the text-image compatibility between the generated images and text. (3) DreamSim is a recently proposed model-based metric to measure perceptual similarity. Similar to image-image CLIPScore, we use DreamSim to compute the perceptual distance between predicted images and ground truth images in a pair-wise manner.
5.2 Main Results
We show the main results of using InterleavedEval to conduct the fine-grained evaluation for various baseline approaches on InterleavedBench in Table 2. The baselines in the upper part are the integrated and open-sourced models while the baselines in the lower part are the pipeline models where the LMMs are proprietary. From Table 2, we observe that: First, the pipeline models consistently outperform the integrated models on all evaluation aspects by a significant margin, where GPT-4o + DALL·E 3 achieves the best performance on helpfulness and the average score of all the aspects. Second, the pipeline models achieve significantly good performance on text quality since Gemini and GPT-4o have strong text generation capabilities. Also, the generated visual prompts are generally coherent with the text content and they are directly fed into the image generation model, so the performance on text-image coherence of pipeline models is also remarkable. Third, we observe that the common errors of integrated models include the output text and/or images being empty, in poor quality, or having severe duplication. This is probably due to their weak instruction-following abilities. Fourth, image coherence is the most challenging aspect for the pipeline models. This is because the image generation model cannot take the images in the input context or previously generated images as conditions. Thus, the generated images do not have strong coherence. We include more qualitative analysis to illustrate these observations in Section 6.
5.3 Human Evaluation
In addition to automatic evaluation, we also conduct an extensive human evaluation to benchmark the baselines and also provide a meta-evaluation on our InterleavedEval and other evaluation metrics by computing the correlation between automatic evaluation scores and human judgments.
Human Evaluation Setup
We adopt the same fine-grained evaluation criteria as InterleavedEval, where for each sample, the annotators need to give a score for each aspect defined in Table 5. The only difference is that, instead of rating on a scale of {0, 1, 2, 3, 4, 5}, we use a scale of {0, 1, 2, 3} for each aspect, where 1, 2, and 3 indicate the quality is bad, fair, and good, respectively. In this way, we can reduce the difficulty of human evaluation and improve its efficiency. Due to the cost of human evaluation, we select four representative baselines to evaluate, i.e., GILL, EMU-2, Gemini1.5 + SDXL, and GPT-4o + DALL·E 3. We include more details on human evaluation setup in Appendix B.1.
Figure 5: Case study. We select the representative examples of the system outputs from GILL, EMU-2, Gemini+SDXL, and GPT-4+DALLE3.
Results
We show the human evaluation results in Table 3. The human evaluation is generally consistent with the automatic evaluation in Table 2. The pipeline models consistently outperform integrated models by a large margin, where GPT-4o+DALL·E 3 also achieves the best performance on helpfulness and the average performance. There’s significant room for improvement in the integrated open-sourced models. We report the Inter Annotator Agreement (IAA) in Table 6 in Appendix B.1.
Correlation Analysis
To validate the effectiveness of our proposed metric, we conduct a correlation analysis by comparing the evaluation results from automatic metrics with our human evaluation results. Since the baseline metrics only predict an overall score for each instance, we use the same set of evaluation scores to compare against the human rating on each aspect separately. For InterleavedEval, we compare evaluation scores with the human rating on corresponding aspects. Since most baselines require a gold reference, we use the context-based subset, where each instance has an associated reference output, to compute the correlation. From Table 4, our InterleavedEval consistently outperforms previous metrics by a significant margin in every aspect. Our metric has a particularly higher correlation on text quality, which is because text quality is relatively easier to evaluate with large language models like GPT-4o Zheng et al. (2023b). Our metric achieves the lowest correlation on perceptual quality. The plausible reason is that GPT-4’s perceptual recognition capability is still not strong enough to accurately detect visual artifacts or unnatural disruptions in the images Fu et al. (2024). We also find that baseline metrics generally achieve poorer correlation, e.g., most metrics except for BERTScore almost do not have any correlation with helpfulness. BERTScore achieves the best correlation on helpfulness among baseline metrics, which indicates that text quality could be a good indicator of whether the overall interleaved content is helpful.
6 Discussions
Figure 6: Radar figure of evaluation results on each evaluation aspect for each task.
Qualitative Analysis
We conduct a qualitative analysis of benchmarked models in Figure 5 and have the following observations: (1) while GILL can generate images with reasonable quality, the generated text and images are typically not coherent with the instruction and context. In the example in the first row, the generated text is totally irrelevant to the task, while the image is also inconsistent with input images. (2) EMU-2 can often generate text that is relevant to the task, but the quality is not good enough. In the example in the second row, it repeatedly says “soak the fabric in water” but does not contain other useful content. Another weakness of EMU-2 is its poor conditional image generation capability, where generated images have obvious visual distortions and could be duplicated with input images. (3) On the other hand, the pipeline models can generally better follow the instructions and generate text and images in higher quality. Nevertheless, they still occasionally have some drawbacks. For Gemini+SDXL, some of the generated images (e.g., the first output image in the second example) still have obvious defects. For GPT-4+DALLE3, the style of generated images can be dramatically different from input images, as DALLE3 is prone to generate images in cartoon or dramatic styles. (4) Maintaining image coherence, i.e., the coherence of style and entities across images, is still very challenging for most models. In the third example, for the pipeline models, the same character has a very different appearance across the images, which makes the content inconsistent. (5) For the instances on the context-free subset, the integrated baselines have significantly worse performance, where they only generate one image with extremely poor quality. We hypothesize the reason to be those models cannot truly understand and follow the instructions. To sum up, our qualitative analysis indicates there is still significant room for improvement in interleaved generation.
Breakdown Results on Each Use Case
We show a detailed breakdown of the average results on all the aspects of each use case. From Figure 6, we observe that (1) for pipeline-based models, image editing and subject-driven generation achieve the lowest results, whereas the models can achieve scores above 4 on other use cases; and (2) integrated models typically achieve low performance on the context-free subset in InterleavedBench. The potential reason is that these models did not specifically fine-turned on the data with text-only inputs, and thus cannot generate interleaved content well.
7 Conclusion
We introduce InterleavedBench, the first benchmark for the evaluation of interleaved text-and-image generation. We also propose InterleavedEval, a strong multi-aspect reference-free evaluation metric based on GPT-4o. With extensive experiments, we first verify that our proposed metric can achieve significantly higher agreement with humans compared with existing metrics. Through the lens of InterleavedEval, we then observed that while the pipeline models based on proprietary LMMs consistently outperform open-source models, interleaved generation is still a challenging task that requires further advancement.
8 Limitation
While our proposed InterleavedBench and InterleavedEval provide a comprehensive evaluation suite for text-and-image interleaved generation, there are still several limitations in our work that we leave for future research. First, while InterleavedEval achieves the best correlation with human judgments among other evaluation metrics, it still does not have a high correlation on certain aspects, such as perceptual quality, image coherence, and text-image coherence. To further improve the evaluation accuracy, we may need to improve the capability of foundation multimodal models such that they are capable of recognizing subtle but critical differences. Second, our work did not extensively address the bias in using GPT-4 for evaluation, which we consider an important topic for future research.
Appendix A More Details on InterleavedBench
Data Filtering Pipeline
To collect the source data from web resources, we first only keep the samples with 3 to 6 images and less than 12 sentences such that the ratio between text and image is balanced. We then apply Llama-8B-Instruct as a text filter to save the data with good text quality. We also apply LPIPS Zhang et al. (2018) to discard the instances with duplicate images.
Manual Data Selection
We apply a manual data selection and instruction annotation process to ensure data quality. We select the instances based on the criteria in Table 5. We also encourage the annotators to select diverse instances.
Instruction Annotation
For each instance, we first ask an annotator to draft an instruction, and then ask another annotator to revise the instruction, until both annotators agree that the instructions are of high quality. The annotators are Ph.D. students with expertise in NLP and multimodal learning areas.
Appendix B More Details on Evaluation
We present the full list of our defined aspects and their definition in Table 5.
B.1 Human Evaluation
More Details on Human Evaluation Setup
We sampled 100 instances from InterleavedBench as a subset for evaluation and ensure its task distribution is the same as the original distribution. In this way, we have 400 data points where each baseline has inference results on 100 instances. For each data point, we have two different annotators who are Ph.D. or master’s students with expertise in NLP or multimodal domains to give ratings independently.
Inter-Annotator Agreement
We show the IAA of our human evaluation in Table 6. While our human evaluation did not achieve significantly high agreement, we argue that the evaluation of interleaved generation is still quite subjective, open-ended, and challenging, even with our carefully designed human evaluation aspects and guidelines.
Appendix C Additional Experiment Results
Aspect Definition
Text Quality Text quality measures how clear, coherent, and error-free the output text is. It considers grammar, spelling, readability, coherence with the instruction and context, and whether it contains duplicate content.
Perceptual Quality Perceptual quality measures how visually convincing, natural, and free from distortions or artifacts a generated image appears. It considers how accurately the image mimics reality without unnatural disruptions in structure, colors, or composition.
Image Coherence Image coherence measures the consistency in style and subject representation across images. This includes textures, color palette, lighting, rendering styles, and maintaining consistent physical attributes, clothing, and behavioral traits. Image coherence also penalizes image duplication, where the output images are too similar, or within the output images themselves.
Text-Image Coherence Text-to-image coherence measure the alignment and integration between textual and visual elements in a pairwise manner, ensuring they work together to convey a unified and cohesive narrative.
Helpfulness Helpfulness measures how well the output text and images follow the task instructions and provide complete information to achieve the task. It also considers whether the outputs follow a reasonable logic flow.
Table 5: The full list of evaluation aspects and their corresponding definitions in InterleavedEval.
Text Quality Perceptual Quality Image Coherence TIC Helpfulness AVG
0.489 0.306 0.320 0.427 0.519 0.412
Table 6: Inter-Annotator Agreement of human evaluation in terms of Cohen’s Kappa score.